To determine the distance between two points A and B, a surveyor chooses a point C that is 355 yards from A and 445 yards from B. If ∠BAC has measure 49°30', approximate the distance between A and B. (Round your answer to the nearest whole number.)
_____yd

Answer by Cromlix(4381)

Hi there,
Using Cosine Rule:
c^2 = a^2 + b^2 - 2abCosC
c^2 = 445^2 + 355^2 - 2 x 445 x 355 x Cos(49.5)
c^2 = 96234.84
c = √96234.84
AB = 310.22 yards (2 decimal places)
Hope this helps:-)